Problems n the problems below, you can use a market risk premium of 3.5% and a tax rate of 40% where none is specified 1. The followin The following is the balance sheet for Ford Motor Company as of December 31, 1994 (in milions: Liabilities Assets $ 11,635 Accounts Payable Cash Receivables Inventory S 19,927 61,469 10,128 91,524 45,586 36,240 Debt due within 1 year Other Current Liabilities Current Liabilities Short-Term Debt Long-Term Debt Equity Total Liabilities 2,721 50,596 36,200 37,490 12,824 Current Assets Fixed Assets Total Assets $137,110 $137,110 The firm had revenues of $154,951 million in 1994 and cost of goods sold of $103,817 million. was consolidated into this balance sheet, would that help you with your explanation? a. Estimate the net working capital b. Estimate the noncash working capital. c. Why is Fords working capital so high? If you were d. Estimate noncash working capital as a percent of revenues. If you were asked to estimate the noncash working capital needs for a new automobile fac- tory that Ford was constructing, would you use this ratio? Why or why not? told that Ford Capital (the financing arm of Ford)
Answer #1
1)
$'million
a. Net Working Capital CA-CL 91524-50596= 40928
b. Non-cash working capital (CA-Cash)- CL (91524-19927) -50596= 21001
c. The Ford's has been keeping a high cash under the working
capital and places a high credit to their customers, so the
working capital of the company is high.
If the financing arm of Ford is consolidated with the Ford Motor
Company's balance sheet, then the cash content of $19927m
belongs to the finance company. Thus, we have to considered
Non-cash working capital for Ford Motors Company.
d. Non-cash working capital (NCWC) as percent of revenues = 21001/154951 = 13.55%
Yes we can use this NCWC for the estimate of new automobile factory
because the Ford is keeping a comfortable and appropriate Net working capital.